1. A 10 kg sack of grain is hoisted to a storage room 50 m above the ground floor of a grain elevator.
a) What is the increase in potential energy of the sack of grain at this height? PE = mgh
1
Expert's answer
2020-12-14T12:14:26-0500

The increase in potential energy is:


"PE = mgh"

where "m = 10kg" is the mass of grain, "h = 50m" is the height, and "g = 9.8m\/s^2" is the gravitational acceleration. Thus, obtain:


"PE = 10kg\cdot 9.8m/s^2\cdot 50m = 4900\space J"

Answer. 4900 J.